# SplatSSC: Decoupled Depth-Guided Gaussian Splatting for Semantic Scene Completion

### v1

@@ Abstract
Monocular 3D Semantic Scene Completion (SSC) is a challenging yet promising task that aims to infer dense geometric and semantic descriptions of a scene from a single image. While recent object-centric paradigms significantly improve efficiency by leveraging flexible 3D Gaussian primitives, they still rely heavily on a large number of randomly initialized primitives, which inevitably leads to 1) inefficient primitive initialization and 2) outlier primitives that introduce erroneous artifacts. In this paper, we propose SplatSSC, a novel framework that resolves these limitations with a depth-guided initialization strategy and a principled Gaussian aggregator. Instead of random initialization, SplatSSC utilizes a dedicated depth branch composed of a Group-wise Multi-scale Fusion (GMF) module, which integrates multi-scale image and depth features to generate a sparse yet representative set of initial Gaussian primitives. To mitigate noise from outlier primitives, we develop the Decoupled Gaussian Aggregator (DGA), which enhances robustness by decomposing geometric and semantic predictions during the Gaussian-to-voxel splatting process. Complemented with a specialized Probability Scale Loss, our method achieves state-of-the-art performance on the Occ-ScanNet dataset, outperforming prior approaches by over 6.3% in IoU and 4.1% in mIoU, while reducing both latency and memory consumption by more than 9.3%. The code will be released upon acceptance.
@@ Introduction
3D scene understanding has garnered significant attention with the rapid evolution of embodied agents and autonomous driving. As a key technology in this domain, 3D occupancy prediction (Tong et al. 2023; Huang et al. 2023; Wei et al. 2023; Tian et al. 2023; Li et al. 2023b; Wang et al. 2024c) and 3D Semantic Scene Completion (SSC) (Cao and de Charette 2022; Miao et al. 2023; Zhang, Zhu, and Du 2023; Li et al. 2023a; Mei et al. 2024) have made remarkable progress. Early and conventional approaches for these tasks predominantly rely on grid-based representations. However, processing dense 3D volumes incurs prohibitive computational and memory costs. To mitigate this limitation, various efficiency-driven strategies have been explored, such as accelerating processing with Bird’s-Eye-View (BEV) projections (Yu et al. 2023; Hou et al. 2024), or leveraging the natural sparsity of scenes with sparse voxels (Tang et al. 2024; Lu et al. 2024; Li et al. 2023a) and points (Shi et al. 2024; Wang et al. 2024a).
@@
Despite the devoted efforts, such methods remain inherently constrained by their discrete and grid-aligned nature, which struggle to model the sparse geometry efficiently. A recent paradigm shift towards object-centric representations, pioneered by GaussianFormer (Huang et al. 2024), has achieved a breakthrough. By utilizing flexible 3D Gaussian primitives (Kerbl et al. 2023) to represent the scene, this approach strikes a new balance between performance and efficiency. Building upon this foundation, subsequent works (Huang et al. 2025) have advanced this field by developing more principled aggregation methods based on Gaussian Mixture Models (GMMs) and adapting the paradigm to indoor scenes for incremental perception (Wu et al. 2025; Zhang et al. 2025; Wang et al. 2025a).
@@
While the object-centric paradigm offers a promising direction, its application in vision-only settings faces a foundational challenge: how to efficiently initialize and reliably supervise 3D primitives using only monocular cues. To ensure complete coverage of the target 3D space without geometric cues, the predominant strategy is to randomly distribute numerous primitives throughout the 3D volume, as shown in Figure 1(a). This leads to two critical, coupled limitations: 1) Inefficient Primitive Initialization. A significant portion of the model’s capacity is inevitably wasted on representing empty or unknown space, making the random distribution strategy inherently redundant. 2) Fragile Aggregation of Outliers. Existing Gaussian-to-voxel splatting strategies (Huang et al. 2024, 2025) lack an effective rejection mechanism to mitigate the impact of outlier primitives. This allows outliers to spurious semantics on distant voxels, creating “floaters” in otherwise empty space.
@@
To this end, we introduce SplatSSC, a novel framework designed to tackle inefficient initialization and fragile aggregation in object-centric SSC. Rather than randomly initializing a large number of primitives, SplatSSC leverages geometric priors to guide the primitive initialization as shown in Figure 1(b), reducing redundancy while maintaining representational capacity. Our methodology begins with a tailored depth branch that generates a high-quality geometric prior through our proposed Group-wise Multi-scale Fusion (GMF) module. GMF integrates multi-scale image features and depth features from a pretrained depth estimator via Group Cross-Attention (GCA) for efficient multi-modal fusion. The resulting geometric priors subsequently guide a lifter to initialize a sparse yet highly targeted set of Gaussian primitives, which are then refined through a standard multi-stage encoder. To address the “floaters” that plague existing aggregators when dealing with sparse outliers, we propose the Decoupled Gaussian Aggregator (DGA), which renders the final semantic grid by completely decomposing semantic and geometry prediction robustly. Furthermore, to ensure stable geometric learning, we design the specialized Probability Scale Loss to apply soft and progressive supervision to the intermediate encoder layers.
@@
3D Semantic Scene Completion (SSC) infers dense geometry and semantics from limited observations, and has recently gained popularity in both indoor and outdoor perception.
Early approaches (Song et al. 2017; Zhang et al. 2019; Wang et al. 2019) primarily focused on indoor scenes using depth-only input, where deep convolutional networks (CNNs) and Truncated Signed Distance Function (TSDF) representations were widely employed. To improve semantic understanding, subsequent methods (Li et al. 2019, 2020; Wang et al. 2023) fuse features from both RGB and depth inputs. In parallel, LiDAR-based SSC approaches (Roldão, de Charette, and Verroust-Blondet 2020; Yan et al. 2021; Yang et al. 2021) have been developed for autonomous driving and also rely on CNN architectures.
A recent trend has shifted towards vision-only methods. MonoScene (Cao and de Charette 2022) pioneered this direction using a dense 2D-to-3D lifting with UNet architecture (Ronneberger, Fischer, and Brox 2015), but this approach suffered from inherent depth ambiguity. To address this, OccDepth (Miao et al. 2023) and ISO (Yu et al. 2024a) introduced depth-aware strategies by leveraging stereo depth and pretrained depth networks, respectively. Concurrently, to tackle the inefficiency of dense voxel processing, VoxFormer (Li et al. 2023a), a two-stage model, proposed a sparse-to-dense Transformer method based on generating proposals from a geometry prior. Subsequent works continue to advance this paradigm (Mei et al. 2024; Yu et al. 2024b; Jiang et al. 2024), focusing on unified pipelines, context-aware modeling, and instance-level reasoning.
While these Transformer-based approaches significantly advance the field, they still operate on grid-aligned voxel queries. Our work diverges by embracing a more flexible object-centric representation, drawing inspiration from recent remarkable works (Wu et al. 2025; Huang et al. 2025).
@@
A recent paradigm shift in occupancy prediction, pioneered by GaussianFormer (Huang et al. 2024), moves beyond grid-aligned queries to object-centric representation using 3D Gaussian primitives. This approach leverages the inherent sparsity of 3D scenes by representing them as a collection of continuous ellipsoids, which are then rendered into a dense semantic grid via an efficient Gaussian-to-voxel splatting mechanism. This marked a significant departure from discrete, voxel-based frameworks (Li et al. 2023a; Tang et al. 2024).
Sequential works (Huang et al. 2025; Zhao et al. 2025) further advanced this paradigm by introducing a principled probabilistic framework via GMMs and incorporating LiDAR-guided initialization to replace random placement. In parallel, EmbodiedOcc (Wu et al. 2025) first adapted this object-centric paradigm to the unique challenges of indoor perception. It focuses on online incremental scene understanding, where confidence refinement is applied to continuously update the Gaussian representation as an agent explores the environment. Following this, RoboOcc and EmbodiedOcc++ (Zhang et al. 2025; Wang et al. 2025a) extended this paradigm through geometry-aware refinement, leveraging opacity cues and planar constraints to enhance stability and structural fidelity.
However, vision-based object-centric approaches widely employ random Gaussian initialization, which introduces significant redundancy, where the majority of primitives are used to represent empty space. In contrast, our method directly tackles this limitation by leveraging a depth prior to generate a compact but more targeted set of primitives.
@@
[…]
@@
The architecture of our approach is illustrated in Figure 2. We first process the input image $\mathcal{I}_{\mathrm{rgb}}$ with an image encoder, composed of a lightweight image backbone EfficientNet (Tan and Le 2019) and FPN (Lin et al. 2017), to extract multi-scale image features $\mathcal{F}_{\mathrm{rgb}}=\{f_{\mathrm{rgb}}^{l}\}_{l=1}^{L}$, where $L$ is the scale number. Simultaneously, a pretrained depth estimation model Depth-Anything (Yang et al. 2024) is employed to produce powerful depth features $\mathcal{F}_{d}$. These two feature streams are then fed into our specialized depth branch, which employs the proposed GMF module to produce the fused depth features $\mathcal{F}_{d}^{\prime}$ and the refined depth map $\mathcal{I}_{d}$. The resulting $\mathcal{F}_{d}^{\prime}$ and $\mathcal{I}_{d}$ are then fed to a lifting module to obtain the initial Gaussian primitives $\mathcal{G}_{o}$ with good geometry prior. Subsequently, $\mathcal{G}_{o}$ is refined by a series of encoder blocks cyclically, following EmbodiedOcc. Given the refined primitives, the 3D semantic voxels are obtained by our DGA $\hat{\mathcal{V}}_{\mathrm{agg}}$. By first leveraging the depth branch to generate a highly compact set of primitives with geometrically grounded initial locations, we tackle the inefficiency inherent in random initialization strategies. Subsequently, our DGA transforms primitives into semantic voxels, overcoming the fragility of prior aggregation methods. This enables our framework to achieve state-of-the-art (SOTA) performance while maintaining high efficiency with significantly fewer primitives.
@@
While recent monocular 3D completion methods (Wu et al. 2025; Yu et al. 2024a) leverage pretrained depth estimators, they tend to utilize depth information as a secondary guiding signal: either refining geometric distributions or informing feature learning. However, this approach neglects the rich latent features generated by depth networks. In contrast, our framework proposes a dual-pronged strategy: we use the depth map as a direct geometric prior, while simultaneously employing the latent depth features as the initial embeddings for 3D primitives. This not only ensures primitives are grounded in both geometry (where) and semantics (what), but necessitates a more advanced fusion mechanism. To fulfill this demand, we design a dedicated depth branch. Inspired by prior works (Ma et al. 2020; Jia et al. 2025), this branch fuses multi-scale image features and depth cues via our GMF mechanism. Specifically, GMF is a Transformer-like block comprising the proposed GCA layer followed by a point-wise FFN (Vaswani et al. 2017). The resulting fused features $\mathcal{F}_{d}^{\prime}$ are then processed by two convolutional layers to produce the refined depth map $\mathcal{I}_{d}$.
@@
The architecture of our GCA module is illustrated in Figure 3. The process begins by sampling features from the input depth features $\mathcal{F}_{d}$ and multi-scale image features $\mathcal{F}_{\mathrm{rgb}}$, using a set of predefined reference points normalized to the [0,1][0,1][ 0 , 1 ] range. This step yields the sampled features, denoted as $\mathcal{F}_{d}^{s}$ and $\mathcal{F}_{\mathrm{rgb}}^{s}=\{f_{\mathrm{rgb}}^{s,l}\}_{l=1}^{L}$ respectively. To balance performance and efficiency, we split these features into $G$ groups along the channel dimension, where each group has a reduced feature dimensionality of $D_{g}=D/G$. The Query $Q_{g}$ is projected from sampled depth features, while the Key $K_{g}^{l}$ and Value $V_{g}^{l}$ are projected from sampled image features at each scale $l$:
@@
where $W_{q}$, $W_{k}$, and $W_{v}$ are linear projection matrices for Query, Key, and Value, respectively. $l\in\{1,...,L\}$ denotes the scale index. Inspired by the efficient design of Deformable Attention (Zhu et al. 2021), we adopt a lightweight linear projection mechanism in place of the standard dot-product attention. To elaborate, the attention scores are computed by feeding the element-wise sum of queries and keys into a shared projection $W_{a}\in\mathbb{R}^{D_{g}\times 1}$
@@
The PSG models the semantic occupancy prediction at a point $\mathbf{x}$ as a two-part process: a geometric occupancy probability $\alpha(\mathbf{x})$ and a conditional semantic expectation $\mathbf{e}(\mathbf{x};\mathcal{G})$:
@@
where $p(\mathbf{x}|G_{i})$ is the Gaussian probability density function, $\alpha(\mathbf{x};G_{i})=\exp(-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu}_{i})^{T}\mathbf{\Sigma}_{i}^{-1}(\mathbf{x}-\boldsymbol{\mu}_{i}))$ is the un-normalized Gaussian kernel. The key flaw in this formulation lies in how the learned opacity $\mathbf{a}_{i}$ is used. While intended to represent a primitive’s existence confidence, it is instead employed as the prior probability in the GMM. The negative consequence of this choice becomes evident when considering an isolated outlier primitive $G_{n}$. For any point $\mathbf{x}^{f}$ in its immediate vicinity, the likelihood $p(\mathbf{x}^{f}|G_{m})$ for all other distant primitives $G_{m\neq n}$ approaches zero. This causes the normalization term in the posterior calculation to be dominated by the outlier itself. Hence, the posterior probability $p(G_{n}|\mathbf{x}^{f})$ collapses to unity, regardless of the effect of the low-confidence prior $\mathbf{a}_{n}$:
@@
Figure 4: Illustration of the proposed DGA. While GF.agg (Huang et al. 2024) and GF2.agg (Huang et al. 2025) wrongly produces the “floaters” from outliers, our DGA remains robust, as the low occupancy probability directly suppresses its erroneous semantic contribution.
@@
Due to the exponential decay of the Gaussian kernel, only primitives in the local vicinity of $\mathbf{x}$ have a meaningful influence. Therefore, we only consider contributions from a neighborhood of relevant Gaussian primitives for efficiency, denoted as $\mathcal{N}(\mathbf{x})$. The occupancy is then modeled as a probabilistic OR operation over this local set. Crucially, each primitive’s influence is modulated by its learned opacity $\mathbf{a}_{i}$, which we interpret as its existence confidence. This explicit use of opacity is a key difference from PSG:
@@
Concurrently, we predict the conditional semantic distribution $e(\mathbf{x})$ under the assumption that the position $\mathbf{x}$ is occupied. This is achieved by using GMM, where we leverage the normalized semantic weights of each Gaussian component. This design decouples the semantic prediction from the opacity parameter $\mathbf{a}_{i}$, forcing the model to rely solely on the geometric proximity and the learned softmax-normalized semantic properties $\tilde{\mathbf{c}}_{i}$ of each primitive. The posterior probability for each semantic class $l$ is then computed as:
$e^{l}(\mathbf{x})=\sum_{i\in\mathcal{N}(\mathbf{x})}p(G_{i}|\mathbf{x})=\frac{\sum_{i\in\mathcal{N}(\mathbf{x})}p(\mathbf{x}|G_{i})\cdot\tilde{\mathbf{c}}_{i}^{l}}{\sum_{j\in\mathcal{N}(\mathbf{x})}p(\mathbf{x}|G_{j})}.$ (11)
@@
Finally, the two decoupled pathways are fused to compute the final probability distribution $\hat{\mathbf{y}}_{x}$ for each 3D position $\mathbf{x}$. The probabilities for each valid semantic class $l$ and the empty class are defined as:
$\left\{\begin{array}[]{l}\hat{\mathbf{y}}_{x}^{l}=\alpha^{\prime}(\mathbf{x})\cdot e^{l}(\mathbf{x})\\ \hat{\mathbf{y}}_{x}^{\mathrm{empty}}=1-\alpha^{\prime}(\mathbf{x})\end{array}\right..$ (14)
This formulation serves as a principled and fully differentiable gating mechanism. A low occupancy probability $\alpha^{\prime}(\mathbf{x})$, often resulting from an outlier primitive, directly suppresses any erroneous semantic prediction $e^{l}(\mathbf{x})$, thus elegantly eliminating “floater” without complex heuristics. We demonstrate this effect in Figure 4.
@@
In this stage, we exclusively train our depth branch to produce a high-quality geometry prior. This module is supervised by a composite depth loss $\mathcal{L}_{\mathrm{d}}$, similar with prior works (Laina et al. 2016; Wang et al. 2025b):
@@
where $\mathcal{L}_{\mathrm{sem}}=\lambda_{5}\mathcal{L}_{\mathrm{focal}}+\lambda_{6}\mathcal{L}_{\mathrm{lovasz}}$ is the primary semantic segmentation loss adopted by EmbodiedOcc. Our loss $\mathcal{L}_{\mathrm{scal}}^{\mathrm{prob}}$ extends the geometry-aware scale loss $\mathcal{L}_{\mathrm{scal}}^{\mathrm{geo}}$ from MonoScene (Cao and de Charette 2022), adapting it to supervise the predicted occupancy probability across all $n$ encoder layers. To account for the progressive refinement across stages, we introduce a linear weighting schedule, which imposes weaker constraints on early-stage predictions and gradually enforces stronger consistency at deeper layers:
@@
where $i$ is the endocer layer index. In our experiments, the loss weights are set as $\lambda_{1}=10$, $\lambda_{2}=20$, $\lambda_{3}=\lambda_{4}=0.5$, $\lambda_{5}=100$, and $\lambda_{6}=2$.
@@
Occ-ScanNet MonoScene $\mathcal{I}_{\mathrm{rgb}}$ 41.60 15.17 44.71 22.41 12.55 26.11 27.03 35.91 28.32 6.57 32.16 19.84 24.62

### v2

@@ Abstract
Monocular 3D Semantic Scene Completion (SSC) is a challenging yet promising task that aims to infer dense geometric and semantic descriptions of a scene from a single image. While recent object-centric paradigms significantly improve efficiency by leveraging flexible 3D Gaussian primitives, they still rely heavily on a large number of randomly initialized primitives, which inevitably leads to 1) inefficient primitive initialization and 2) outlier primitives that introduce erroneous artifacts. In this paper, we propose SplatSSC, a novel framework that resolves these limitations with a depth-guided initialization strategy and a principled Gaussian aggregator. Instead of random initialization, SplatSSC utilizes a dedicated depth branch composed of a Group-wise Multi-scale Fusion (GMF) module, which integrates multi-scale image and depth features to generate a sparse yet representative set of initial Gaussian primitives. To mitigate noise from outlier primitives, we develop the Decoupled Gaussian Aggregator (DGA), which enhances robustness by decomposing geometric and semantic predictions during the Gaussian-to-voxel splatting process. Complemented with a specialized Probability Scale Loss, our method achieves state-of-the-art performance on the Occ-ScanNet dataset, outperforming prior approaches by over $6.3\%$ in IoU and $4.1\%$ in mIoU, while reducing both latency and memory cost by more than $9.3\%$.
Codehttps://github.com/Made-Gpt/SplatSSC
@@ Introduction
3D scene understanding has garnered significant attention with the rapid evolution of embodied agents (deng2025mcnslammultiagentcollaborativeneural; deng2025mne; deng2024plgslam) and autonomous driving (li2025flagdroneacing; yi2025embodiedgame; fusic2024improvedrrt; zheng2025robustefficient; zammit2023realtimeuav). As a key technology in this domain, 3D occupancy prediction (tong2023scene; huang2023tri; wei2023surroundocc; tian2023occ3d; wang2024panoocc) and 3D Semantic Scene Completion (SSC) (cao2024monoscene; miao2023occdepth; zhang2023occformer; li2023voxformer; mei2024camera) have made remarkable progress. Early and conventional approaches for these tasks predominantly rely on grid-based representations. However, processing dense 3D volumes incurs prohibitive computational and memory costs. To mitigate this limitation, various efficiency-driven strategies have been explored, such as accelerating processing with Bird’s-Eye-View (BEV) projections (yu2023flashocc; hou2024fastocc), or leveraging the natural sparsity of scenes with sparse voxels (tang2024sparseocc; li2023voxformer) and points (shi2024occupancy; wang2024opus).
@@
Despite the devoted efforts, such methods remain inherently constrained by their discrete and grid-aligned nature, which struggle to model the sparse geometry efficiently. A recent paradigm shift towards object-centric representations, pioneered by GaussianFormer (huang2024gaussianformer), has achieved a breakthrough. By utilizing flexible 3D Gaussian primitives (kerbl20233d) to represent the scene, this approach strikes a new balance between performance and efficiency. Building upon this foundation, subsequent works (huang2025gaussianformer2) have advanced this field by developing more principled aggregation methods based on Gaussian Mixture Models (GMMs) and adapting the paradigm to indoor scenes for incremental perception (wu2024embodiedocc; zhang2025roboocc; wang2025embodiedoccplusplus).
@@
While the object-centric paradigm offers a promising direction, its application in vision-only settings faces a foundational challenge: how to efficiently initialize and reliably supervise 3D primitives using only monocular cues. To ensure complete coverage of the target 3D space without geometric cues, the predominant strategy is to randomly distribute numerous primitives throughout the 3D volume, as shown in Figure 1(a). This leads to two critical, coupled limitations: 1) Inefficient Primitive Initialization. A significant portion of the model’s capacity is inevitably wasted on representing empty or unknown space, making the random distribution strategy inherently redundant. 2) Fragile Aggregation of Outliers. Existing Gaussian-to-voxel splatting strategies (huang2024gaussianformer; huang2025gaussianformer2) lack an effective rejection mechanism to mitigate the impact of outlier primitives. This allows outliers to spurious semantics on distant voxels, creating “floaters” in otherwise empty space.
@@
To this end, we introduce SplatSSC, a novel framework designed to tackle inefficient initialization and fragile aggregation in object-centric SSC. Rather than starting with a large set of random primitives, SplatSSC leverages geometric priors to guide the primitive initialization, as shown in Figure 1(b), reducing redundancy while maintaining representational capacity. We generate this prior using a tailored depth branch, equipped with our proposed Group-wise Multi-scale Fusion (GMF) module. GMF integrates multi-scale image features and depth features from a pretrained depth estimator via Group Cross-Attention (GCA) for efficient multi-modal fusion. The resulting geometric priors subsequently guide a lifter to initialize a sparse yet targeted set of Gaussian primitives, which are then refined through a standard multi-stage encoder. To address the “floaters” that plague existing aggregators when dealing with sparse outliers, we propose the Decoupled Gaussian Aggregator (DGA), which renders the final semantic grid by completely decomposing semantic and geometry prediction robustly. Furthermore, to ensure stable geometric learning, we design the specialized Probability Scale Loss to apply soft and progressive supervision to the intermediate encoder layers.
@@
3D Semantic Scene Completion (SSC) infers dense geometry and semantics from partial observations. Early approaches (song2017sscnet; zhang2019cascaded; wang2019forknet) primarily focused on indoor scenes using depth-only input, where deep convolutional networks (CNNs) and Truncated Signed Distance Function (TSDF) representations were widely employed. To improve semantic understanding, subsequent methods (li2019rgbd; li2020anisotropic; wang2023semantic) fuse features from both RGB and depth inputs. In parallel, LiDAR-based SSC approaches (roldao2020lmscnet; yan2021sparse; yang2021semantic) have been developed for autonomous driving and also rely on CNN architectures.
A recent trend has shifted towards vision-only methods. MonoScene (cao2024monoscene) pioneered this direction using a dense 2D-to-3D lifting with UNet architecture (ronneberger2015u), but this approach suffered from inherent depth ambiguity. To address this, OccDepth (miao2023occdepth) and ISO (yu2024monocular) introduced depth-aware strategies by leveraging stereo depth and pretrained depth networks, respectively. Concurrently, to tackle the inefficiency of dense voxel processing, VoxFormer (li2023voxformer), a two-stage model, proposed a sparse-to-dense Transformer method based on generating proposals from a geometry prior. Subsequent works continue to advance this paradigm (mei2024camera; zhu2024CGFormer; jiang2024symphonize), focusing on unified pipelines, context-aware modeling, and instance-level reasoning.
While these Transformers improve accuracy, they remain bound to grid-aligned voxel queries. Our work takes a different route through a flexible object-centric formulation inspired by (wu2024embodiedocc; huang2025gaussianformer2).
@@
A recent paradigm shift in occupancy prediction, pioneered by GaussianFormer (huang2024gaussianformer), moves beyond grid-aligned queries to object-centric representation using 3D Gaussian primitives. This approach leverages the inherent sparsity of 3D scenes by representing them as a collection of continuous ellipsoids, which are then rendered into a dense semantic grid via an efficient Gaussian-to-voxel splatting mechanism. This marked a significant departure from discrete, voxel-based frameworks (li2023voxformer; tang2024sparseocc).
Sequential works (huang2025gaussianformer2; zhao2025gaussianformer3d) further advanced this paradigm by introducing a principled probabilistic framework via GMMs and incorporating LiDAR-guided initialization to replace random placement. In parallel, EmbodiedOcc (wu2024embodiedocc) first adapted this object-centric paradigm to the unique challenges of indoor perception. It focuses on online incremental scene understanding, where confidence refinement is applied to continuously update the Gaussian representation as an agent explores the environment. Following this, RoboOcc and EmbodiedOcc++ (zhang2025roboocc; wang2025embodiedoccplusplus) extended this paradigm through geometry-aware refinement, leveraging opacity cues and planar constraints to enhance stability and structural fidelity.
However, object-centric approaches widely employ random Gaussian primitive initialization, which introduces significant redundancy, as most primitives are used to represent empty space. In contrast, our method directly tackles this problem by leveraging a depth prior to generate a compact but more targeted set of primitives.
@@
where $\mathrm{q2r}(\cdot)$ converts a quaternion into a rotation matrix and $\mathrm{diag}(\cdot)$ forms a diagonal scaling matrix.
@@
The architecture of our approach is illustrated in Figure 2. We first process the input image $\mathcal{I}_{\mathrm{rgb}}$ with an image encoder, composed of a lightweight image backbone EfficientNet (tan2019efficientnet) and FPN (lin2017feature), to extract multi-scale image features $\mathcal{F}_{\mathrm{rgb}}=\{f_{\mathrm{rgb}}^{l}\}_{l=1}^{L}$, where $L$ is the scale number. Simultaneously, a pretrained depth estimation model Depth-Anything (yang2024depth) is employed to produce powerful depth features $\mathcal{F}_{d}$. These two feature streams are then fed into our specialized depth branch, which employs the proposed GMF module to produce the fused depth features $\mathcal{F}_{d}^{\prime}$ and the refined depth map $\mathcal{I}_{d}$. The resulting $\mathcal{F}_{d}^{\prime}$ and $\mathcal{I}_{d}$ are then fed to a lifting module to obtain the initial Gaussian primitives $\mathcal{G}_{o}$ with good geometry prior. Subsequently, $\mathcal{G}_{o}$ is refined by a series of encoder blocks cyclically, following EmbodiedOcc. Given the refined primitives, the 3D semantic voxels are obtained by our DGA $\hat{\mathcal{V}}_{\mathrm{agg}}$. By first leveraging the depth branch to generate a highly compact set of primitives with geometrically grounded initial locations, we tackle the inefficiency inherent in random initialization strategies. Subsequently, our DGA transforms primitives into semantic voxels, overcoming the fragility of prior aggregation methods. This enables our framework to achieve state-of-the-art (SOTA) performance while maintaining high efficiency with significantly fewer primitives.
@@
While recent monocular 3D completion methods (wu2024embodiedocc; yu2024monocular) leverage pretrained depth estimators, they tend to utilize depth information as a secondary guiding signal: either refining geometric distributions or informing feature learning. However, this approach neglects the rich latent features generated by depth networks. In contrast, our framework proposes a dual-pronged strategy: we use the depth map as a direct geometric prior, while simultaneously employing the latent depth features as the initial embeddings for 3D primitives. This not only ensures primitives are grounded in both geometry (where) and semantics (what), but necessitates a more advanced fusion mechanism. To fulfill this demand, we design a dedicated depth branch. Inspired by prior works (ma2020auto; jia2025gated), this branch fuses multi-scale image features and depth cues via our GMF mechanism. Specifically, GMF is a Transformer-like block comprising the proposed GCA layer followed by a point-wise FFN (ashish2017transformer). The resulting fused features $\mathcal{F}_{d}^{\prime}$ are then processed by two convolutional layers to produce the refined depth map $\mathcal{I}_{d}$.
@@
The architecture of our GCA module is illustrated in Figure 3. The process begins by sampling features from the input depth features $\mathcal{F}_{d}$ and multi-scale image features $\mathcal{F}_{\mathrm{rgb}}$, using a set of predefined reference points normalized to the $[0,1]$ range. This step yields the sampled features, denoted as $\mathcal{F}_{d}^{s}$ and $\mathcal{F}_{\mathrm{rgb}}^{s}=\{f_{\mathrm{rgb}}^{s,l}\}_{l=1}^{L}$ respectively. To balance performance and efficiency, we split these features into $G$ groups along the channel dimension, where each group has a reduced feature dimensionality of $D_{g}=D/G$. The Query $Q_{g}$ is projected from sampled depth features, while the Key $K_{g}^{l}$ and Value $V_{g}^{l}$ are projected from sampled image features at each scale $l$:
@@
where $W_{q}$, $W_{k}$, and $W_{v}$ are linear projection matrices for Query, Key, and Value, respectively. $l\in\{1,...,L\}$ denotes the scale index. Inspired by the efficient design of Deformable Attention (zhu2020deformable), we adopt a lightweight linear projection mechanism in place of the standard dot-product attention. To elaborate, the attention scores are computed by feeding the element-wise sum of queries and keys into a shared projection $W_{a}\in\mathbb{R}^{D_{g}\times 1}$
@@
The PGS models the semantic occupancy prediction at a point $\mathbf{x}$ as a two-part process: a geometric occupancy probability $\alpha(\mathbf{x})$ and a conditional semantic expectation $\mathbf{e}(\mathbf{x};\mathcal{G})$:
@@
where $p(\mathbf{x}\mid G_{i})$ denotes the conditional Gaussian probability of point $\mathbf{x}$ under the $i$th primitive, and $p(G_{i}\mid\mathbf{x})$ represents the normalized posterior of selecting $G_{i}$ for $\mathbf{x}$ under a uniform prior. $\alpha(\mathbf{x};G_{i})=\exp(-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu}_{i})^{T}\mathbf{\Sigma}_{i}^{-1}(\mathbf{x}-\boldsymbol{\mu}_{i}))$ is the un-normalized Gaussian kernel. The key flaw in this formulation lies in how the learned opacity $\mathbf{a}_{i}$ is used. While intended to represent a primitive’s existence confidence, it is instead employed as the prior probability in the GMM. The negative consequence of this choice becomes evident when considering an isolated outlier primitive $G_{n}$. For any point $\mathbf{x}^{f}$ in its immediate vicinity, the likelihood $p(\mathbf{x}^{f}|G_{m})$ for all other distant primitives $G_{m\neq n}$ approaches zero. This causes the normalization term in the posterior calculation to be dominated by the outlier itself. Hence, the posterior probability $p(G_{n}|\mathbf{x}^{f})$ collapses to unity, regardless of the effect of the low-confidence prior $\mathbf{a}_{n}$:
@@
Figure 4: Illustration of the proposed DGA. While GF.agg (huang2024gaussianformer) and GF2.agg (huang2025gaussianformer2) wrongly produces the “floaters” from outliers, our DGA remains robust, as the low occupancy probability directly suppresses its erroneous semantic contribution.
@@
Due to the exponential decay of the Gaussian kernel, only primitives in the local vicinity of $\mathbf{x}$ have a meaningful influence. Therefore, we only consider contributions from a neighborhood of relevant Gaussian primitives for efficiency, denoted as $\mathcal{N}(\mathbf{x})$. The occupancy is then modeled as a probabilistic OR operation over this local set. Crucially, each primitive’s influence is modulated by its learned opacity $\mathbf{a}_{i}$, which we interpret as its existence confidence. This explicit use of opacity is a key difference from PGS:
@@
Concurrently, we predict the conditional semantic distribution $e(\mathbf{x})$ under the assumption that the position $\mathbf{x}$ is occupied. This is achieved by using GMM, where we leverage the normalized semantic weights of each Gaussian component. This design decouples the semantic prediction from the opacity parameter $\mathbf{a}_{i}$, forcing the model to rely solely on the geometric proximity and the learned softmax-normalized semantic properties $\tilde{\mathbf{c}}_{i}$ of each primitive. The posterior probability for each semantic class $k$ is then computed as:
$e^{k}(\mathbf{x})=\sum_{i\in\mathcal{N}(\mathbf{x})}p(G_{i}|\mathbf{x})=\frac{\sum_{i\in\mathcal{N}(\mathbf{x})}p(\mathbf{x}|G_{i})\cdot\tilde{\mathbf{c}}_{i}^{k}}{\sum_{j\in\mathcal{N}(\mathbf{x})}p(\mathbf{x}|G_{j})}.$ (11)
@@
Finally, the two decoupled pathways are fused to compute the final probability distribution $\hat{\mathbf{y}}_{x}$ for each 3D position $\mathbf{x}$. The probabilities for each valid semantic class $k$ and the empty class are defined as:
$\left\{\begin{array}[]{l}\hat{\mathbf{y}}_{x}^{k}=\alpha^{\prime}(\mathbf{x})\cdot e^{k}(\mathbf{x})\\ \hat{\mathbf{y}}_{x}^{\mathrm{empty}}=1-\alpha^{\prime}(\mathbf{x})\end{array}\right.$. (14)
This formulation serves as a principled and fully differentiable gating mechanism. A low occupancy probability $\alpha^{\prime}(\mathbf{x})$, often resulting from an outlier primitive, directly suppresses any erroneous semantic prediction $e^{k}(\mathbf{x})$, thus elegantly eliminating “floaters” without complex heuristics. We demonstrate this effect in Figure 4.
@@
In this stage, we exclusively train our depth branch to produce a high-quality geometry prior. This module is supervised by a composite depth loss $\mathcal{L}_{\mathrm{d}}$, similar with prior works (laina2016deeper; wang2025vggt):
@@
where $\mathcal{L}_{\mathrm{sem}}=\lambda_{5}\mathcal{L}_{\mathrm{focal}}+\lambda_{6}\mathcal{L}_{\mathrm{lovasz}}$ is the primary semantic segmentation loss adopted by EmbodiedOcc. Our loss $\mathcal{L}_{\mathrm{scal}}^{\mathrm{prob}}$ extends the geometry-aware scale loss $\mathcal{L}_{\mathrm{scal}}^{\mathrm{geo}}$ from MonoScene (cao2024monoscene), adapting it to supervise the predicted occupancy probability across all $n$ encoder layers. To account for the progressive refinement across stages, we introduce a linear weighting schedule, which imposes weaker constraints on early-stage predictions and gradually enforces stronger consistency at deeper layers:
@@
where $i$ is the layer index. The loss weights are set as $\lambda_{1}=10$, $\lambda_{2}=20$, $\lambda_{3}=\lambda_{4}=0.5$, $\lambda_{5}=100$, and $\lambda_{6}=2$.
@@
Occ-ScanNet TPVFormer $\mathcal{I}_{\mathrm{rgb}}$ 33.39 6.96 32.97 14.41 9.10 24.01 41.49 45.44 28.61 10.66 35.37 25.31 24.94
GaussianFormer $\mathcal{I}_{\mathrm{rgb}}$ 40.91 20.70 42.00 23.40 17.40 27.00 44.30 44.80 32.70 15.30 36.70 25.00 29.93
MonoScene $\mathcal{I}_{\mathrm{rgb}}$ 41.60 15.17 44.71 22.41 12.55 26.11 27.03 35.91 28.32 6.57 32.16 19.84 24.62